From a group of 10 boys and 12 girls, a committee of 4 students is chosen at random. How many four-person committees containing at least 1 girl are possible?
Answer by sudhanshu_kmr(1152)


here we have to find no. of groups in which at least one girl

i.e from total no. of groups discard those groups in which only boys.


no. of ways to selecting 4 from total 22 = 22C4

no. of ways to selecting 4 boys from total 10 boys = 10C4


total possible ways = 22C4 - 10C4 = 7315 - 210 = 7105


answer is: 7105


It is possible that some typing mistake in solution of a problem, if any please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......